INFRARED-ULTRAVIOLET DOUBLE RESONANCE SPECTROSCOPY OF ACETYLENE: DYNAMICAL SYMMETRY BREAKING IN THE 4ν_{CH} ROVIBRATIONAL MANIFOLD AT 12700 cm^{−1}

MARK A. PAYNE, \underline{ANGELA P. MILCE}, MICHAEL J. FROST and BRIAN J. ORR, School of Chemistry and Centre for Lasers and Applications, Macquarie University, Sydney, NSW, Australia 2109.

Time-resolved fluorescence-detected infrared-ultraviolet optical double resonance (IR-UV DR) experiments have been performed with a Raman-shifted dye laser preparing gas-phase C_{2}H_{2} molecules in the ”4ν_{CH}” region followed by laserinduced fluorescence (LIF) probing in the S_{1} ← S_{0} vibronic band system. Unusual symmetry-breaking energy transfer has been observed, induced (at least in part) by collisions.^{a} This takes the form of odd-numbered changes of the rotational quantum number J, despite the fact that intramolecular transfer between the ortho and para nuclear-spin modifications of such a molecule is usually forbidden.

The selection rules for IR-UV DR spectra determine that the final vibronic levels should be of gerade symmetry. UV spectra are obtained by IR excitation of particular rotational levels of the (10300)^{0} vibrational eigenstate that terminate in ungerade vibronic levels. Odd-numbered changes of J are also observed in these spectra.

We have verified that the mechanism is an intramolecular process. One possibility is collision-induced dynamical symmetry breaking involving transitions between ”+” and ”-” local-mode states which spoil the g / u symmetry. We also consider whether a / s nuclear-spin interchange symmetry can be involved.

aM. A. Payne, A. P. Milce, M. J. Frost and B. J. Orr, Chem. Phys. Lett. \underline{265}, 244 (1997).